OpenStudy (anonymous):

Whats the derivative of 4pi(r)^3 using the first principles? Show steps. Im asking because I got 4pi(r)^2 and I know its 12pi(r)^2

OpenStudy (anonymous):

first principles means limit definition?

OpenStudy (anonymous):

nvm I got it :P

OpenStudy (anonymous):

\[\lim_{h\to 0}\frac{4\pi(r+h)^2-4\pi r^2}{h}\]

OpenStudy (anonymous):

Yes, in my textbook, its called first principles